Find the x-intercepts of the graph of the equations:
y = x^2 - 3x + 2

y = x^2 - 1

Answer by ewatrrr(24785)


Hi,
x-intercepts are dtyermined when y = 0
x^2 - 3x + 2 = 0
factoring
(x-2)(x-1) = 0 Note:SUM of the inner product(-2x) and the outer product(-x) = -3x
(x-2)= 0 x = 2 x -Intercept Pt(2,0)
(x-1) = 0 x = 1 x -Intercept Pt(1,0)

x^2-1 = 0
x^2 = 1
x = ± 1 x -Intercepts: Pt(1,0)and Pt(-1,1)